# version 1.0

### v6

@@ 5.6 Exercises
5.2. Derive the solution to the infinite-horizon LQR with zero mean Gaussian noise.
5.3. Extend the proof of Theorem 5.2 to positive semidefinite $Q,R$, rather than positive definite matrices.
5.4. Prove Theorem 5.4 as a dynamic two-player zero-sum game, where a unique saddle-point solution exists. Hint: you can formulate this problem as a min-max optimization problem within a finite horizon.
5.5. Prove that the function $f(\mathbf{x})=\mathbf{x}^{\top}M\mathbf{x}$ is convex if and only if $M\succeq 0$ is positive semidefinite.
@@ 6.3.1 Linear Policies
A linear controller $K\in{\mathbb{R}}^{d_{x}\times d_{u}}$ is a linear operator that maps the state to the control. We denote by $\Pi_{\gamma}^{L}$ the set of all linear controllers that are $\gamma$-stabilizing according to definition 2.4 and are bounded in Frobenius norm by $\|K\|\leq\gamma$.
@@ 6.3.2 Disturbance Action Controllers
Any DAC policy in $\Pi_{h,\gamma}^{D}$ is $\gamma^{2}$-stabilizing.
@@
For convenience, denote by $M_{0}=I$, and the fictitious noise term $\tilde{\mathbf{w}}_{t}=\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i}$. First, due to the bound on $M_{1:h}$ and $\mathbf{w}_{t}$, we have that
$\|\tilde{\mathbf{w}}_{t}\|=\|\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i}\|\leq\sum_{i=0}^{h}\|M_{i}\|\|\mathbf{w}_{t-i}\|\leq\gamma,$
@@
$=A\mathbf{x}_{t}+B(K\mathbf{x}_{t}+\sum_{i=1}^{h}M_{i}\mathbf{w}_{t-i})+\mathbf{w}_{t}$
$=(A+BK)\mathbf{x}_{t}+B\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i}$ $\text{using }M_{0}=I$
$=\sum_{j=0}^{t}\left((A+BK)^{j}B\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i-j}\right)$
$=\sum_{j=0}^{t-1}\left((A+BK)^{j}B\tilde{\mathbf{w}}_{t-j}\right).$
We can thus view the states $x_{t}$ as generated by a $\gamma$-stabilizing policy over noises that are of magnitude $\gamma$, and hence bounded by $\gamma^{2}$. ∎
@@
$\leq\left\|\sum_{i=0}^{t}A^{i}B\right\|\max_{i\leq t}\|\mathbf{u}_{t}-\tilde{\mathbf{u}}_{i}\|\leq\frac{\kappa}{\delta}\varepsilon,$
@@ 6.4 Bibliographic Remarks
Regret minimization, online learning, and online convex optimization have a rich history and have been studied for decades in the machine learning and game theory literature. The reader is referred to the comprehensive texts of [CBL06, Haz16] for a detailed treatment.
Linear state-feedback policies are a mainstay of control theory in both applications, due to their simplicity, and in theory, due to their optimality in deterministic and stochastic settings, as demonstrated in [Kal60]. The methods for computing such policies are often based on dynamic programming. Often motivated by the need to impose constraints beyond optimality, such as robustness, various other policy parameterizations were introduced. An early example is the Youla parameterization, sometimes called Q-parameterization [YJB76]. Later, the introduction of system-level synthesis (SLS) [DMW+17, ADLM19], in a similar vein, proved useful for distributed control, especially given its emphasis on state-space representations. For a comparison of these and additional ones such as input-output parameterizations [FZPK19], see [ZFP+20, ZZLL23].
The Disturbance Action Control parameterization, introduced in [ABH+19], is similar to system-level synthesis. Like SLS, it is a convex parameterization that is amenable to the tools of convex optimization. However, unlike SLS, it is a more compact representation that has logarithmically many parameters in its natural representation in terms of the desired approximation guarantee against linear policies. This makes it better suited for machine learning.
The study of regret in the context of control has roots in the work of [AYS11, AYBK14]. These works, as well as many that followed [DMM+18, MTR19, CHK+18, AHS19, SF20], study regret in a stochastic setting, where the perturbations are stochastic. In such settings, optimal control is well defined. The full power of regret analysis for control, involving both adversarial perturbations, adversarial loss functions, and computational considerations, was initiated in [ABH+19]. The notion of regret as applied to policy classes was implicit in the work of [EDKM04]. The regret metric when applied to policies is different from the counterfactual notion of policy regret defined by [ADT12].
@@ 7.4 Bibliographic Remarks
The bulk of the material in this chapter is from [ABH+19]. The work of [ABH+19] defined the general online nonstochastic setting in which the disturbances and the cost functions are adversarially chosen, and where the cost functions are arbitrary convex functions.
The follow-up work by [AHS19] achieves a logarithmic pseudo-regret for strongly convex, adversarially selected losses, and well-conditioned stochastic noise. Logarithmic regret algorithms for the full setting of online nonstochastic control with strongly convex costs were obtained in [Sim20, FS20, CCK20].
Online nonstochastic control with known dynamics was further extended in several directions. The task of control with bandit feedback for the costs was considered in [GHH20]. More refined regret metrics applicable to changing dynamics were studied in [GHM20, MGSH21, ZWZ22, Sin22]. More extensions suitable for changing environments were studied in [BW22]. Online nonstochastic control of population dynamics was studied in [GHL+24].
[LDL21] studies algorithms that balance regret minimization while adhering to constraints. An application of nonstochastic control in iterative planning in the presence of perturbations is described in [AHMS21]. Various open-source software packages, such as [GHM20], provide benchmarked implementations of many such algorithms.
An alternate performance metric for online control, which is also nonstochastic, is the competitive ratio. This objective tracks the cost of the online controller and compares it to the best policy in hindsight. Control methods that minimize the competitive ratio were initiated in [GW19]. This approach was extended in a series of papers [GLSW19, SLC+20, GH21]. Thus far, the success of this approach was limited to fully observed linear dynamical systems and further to quadratic costs when achieving the optimal competitive ratio. [GASH22] show that regret minimizers in the non-stochastic setting automatically achieve a near-optimal competitive ratio.
A precursor to online nonstochastic control is a line of work, beginning with [AYBK14], focused on the control of stochastic systems under sequentially revealed adversarial costs functions. See [CK20, CCK22] for further developments.
@@ 8.6 Bibliographic Remarks
In contrast, much of the work at the intersection of machine learning and control is focussed on precise non-asymptotic bounds on sample complexity or regret. [Fie97] which studies the discounted LQR problem under stochastic noise is the earliest example that provides a precise sample complexity bound. Similarly, the early work of [AYS11] uses an optimism-based approach to achieve optimal regret in terms of dependence on horizon. This regret bound was later improved for sparse systems in [IJR12]. [DMM+18, DMM+20] established regret and sample complexity results for the general stochastic LQR formulation with a polynomial dependence on natural system parameters. [FGKM18] studies the efficacy of first-order optimization-based approaches in synthesizing controllers for the LQR problem. The work of [MTR19] establishes that simple explore-then-commit strategies achieve the optimal dependence of regret on horizon, with [SF20] proving that such approaches are optimal in terms of their dependence on dimension of the system, as well. [CKM19] also gave a first optimal-in-horizon polynomial regret result simultaneously with the former work, albeit using an optimism-based approach. The result of [PH20] gives a rate-optimal regret bound for unknown systems with convex costs.
@@ 9.1 Learning in Dynamical Systems
For the remainder of this chapter, we consider a partially observable linear time-invariant dynamical system described by the following equations. The extension of the material in this chapter to time-varying linear dynamical systems is left as exercise 9.7.
$\mathbf{x}_{t+1}$ $=A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t}$
@@ 9.2 Online Learning of Dynamical Systems
In an online learning setting, the learner sequentially observes the data and makes predictions in real time. Formally, the process unfolds as follows:
  1. 1.
    At each time step $t$, the learner observes the control input $\mathbf{u}_{t-1}$.
  2. 2.
    The learner makes a prediction $\hat{\mathbf{y}}_{t}$ for the next output.
  3. 3.
    The true output $\mathbf{y}_{t}$ is revealed.
  4. 4.
    The learner incurs a loss based on the difference between $\mathbf{y}_{t}$ and $\hat{\mathbf{y}}_{t}$, using a predefined loss function $\ell(\hat{\mathbf{y}}_{t},\mathbf{y}_{t})$, such as squared error:
    $\ell(\hat{\mathbf{y}}_{t},\mathbf{y}_{t})=\|\hat{\mathbf{y}}_{t}-\mathbf{y}_{t}\|^{2}.$
The goal is to minimize the total loss over time. However, in adversarial or unpredictable settings, direct loss minimization may be ill-posed. Instead, we consider a regret-based formulation, where the learner’s performance is measured relative to the best possible predictor from a predefined class of functions.
More precisely, let $\Pi$ be a family of prediction rules mapping past observations to predictions. The regret of an online learning algorithm $A$ with respect to $\Pi$ is given by:
$\text{Regret}_{T}(A,\Pi)=\sum_{t=1}^{T}\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|^{2}-\min_{\pi\in\Pi}\sum_{t=1}^{T}\|\mathbf{y}_{t}-\hat{\mathbf{y}}^{\pi}_{t}\|^{2}$
where $\hat{\mathbf{y}}^{\pi}_{t}$ denotes the prediction made at time $t$ by the best function $\pi\in\Pi$ in hindsight.
This regret-based framework allows us to compare different classes of predictors and design learning algorithms that efficiently adapt to unknown dynamical systems.
@@ 9.3 Classes of Prediction Rules
We say that a class of predictors $\Pi_{1}$ $\varepsilon$-approximates class $\Pi_{2}$ if the following holds. For every sequence of $T$ inputs and outputs $\{\mathbf{u}_{1:T},\mathbf{y}_{1:T}\}$, and every $\pi_{2}\in\Pi_{2}$, there exists $\pi_{1}\in\Pi_{1}$ such that,
$\sum_{t=1}^{T}\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|\leq T\varepsilon,$
where $\mathbf{y}_{t}^{\pi_{1}}=\pi_{1}(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1})\ ,\ \mathbf{y}_{t}^{\pi_{2}}=\pi_{2}(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1})$
@@
Let predictor class $\Pi_{1}$ $\varepsilon$-approximate class $\Pi_{2}$. Then for any prediction algorithm $\mathcal{A}$} we have
@@ 9.3.1 Linear Predictors
The simplest prediction rule is a linear function of observations, inputs, or both. More precisely,
Definition 9.3 (Linear Policies).
A linear controller parametrized by matrices $\{M_{1,i},M_{2,j}\}\in{\mathbb{R}}^{h\times d_{y}\times d_{y}+k\times d_{y}\times d_{u}}$ is a linear operator mapping state to control as
@@
We denote by $\Pi_{h,k,\kappa}^{L}$ the set of all linear controllers bounded as $\sum_{i=1}^{h}\|M_{1,i}\|+\sum_{j=1}^{k}\|M_{2,j}\|\leq\kappa$.
@@ 9.3.2 Linear Dynamical Predictors
An intuitive class of prediction rules that makes use of the special structure of linear dynamical systems is that of open-loop linear dynamical predictors, which we now define.
The motivation for this class is the case that the system transformations are known and the system is fully observed, namely $\mathbf{C}=I$, the identity matrix. In this situation, the optimal prediction can be calculated using the system evolution equation $\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t}$. It can be seen that this prediction rule is optimal for any zero-mean noise or perturbation model. However, this optimality is only true if the state is fully observable. In the next chapter, we consider optimal predictors for zero-mean i.i.d. noise in the more general partially observed case.
@@
$=CA\mathbf{x}_{t-1}+CB\mathbf{u}_{t-1}$
$\vdots$
$=\sum_{i=1:t-1}CA^{i-1}B\mathbf{u}_{t-i}+CA^{t}\mathbf{x}_{0}.$
Thus, the operator $\pi_{\{A,B,C,\mathbf{x}_{0}\}}$ takes a history of states and controls and gives the next output of the system accurately. We denote by $\prod^{\star}$ the set of all such predictors that in the case of zero noise give a perfectly accurate prediction.
If the spectral radius of $A$ is larger than one, then the system is unstable and the magnitude of $\mathbf{y}_{t}$ can grow arbitrarily large, as discussed in section 4.2. Thus, for the rest of this section, assume that $A$ is $\gamma$-stable.
With a decaying system, we can limit the memory of the predictor and still maintain accuracy. Thus, it makes sense to consider a more restricted class of predictors that have bounded memory, that is,
$\pi_{\{A,B,C,h\}}=\sum_{i=1}^{h}CA^{i-1}B\mathbf{u}_{t-i},$
where as a matter of convention we say $\mathbf{u}_{t}=0$ for all negative $t$. Denote by $\prod_{h}^{\star}$ the class of all such predictors.
Clearly, for a large $h$ this class of predictors well approximates the true observations $\mathbf{y}_{t}$ if the system is activated without any noise. We next formalize what this means.

### v8

@@ 5.6 Exercises
5.2. Formulate the infinite-horizon LQR with i.i.d. zero-mean Gaussian noise as an average-cost control problem, and derive the corresponding Bellman equation and optimal linear policy.
5.3. Extend the proof of Theorem 5.2 to the case $Q\succeq 0$ and $R\succ 0$.
5.4. Prove Theorem 5.3 by formulating the robust control problem above as a finite-horizon dynamic two-player zero-sum game.
5.5. Prove that the function $f(\mathbf{x})=\mathbf{x}^{\top}M\mathbf{x}$ is convex if and only if the symmetric part $\frac{1}{2}(M+M^{\top})$ is positive semidefinite.
@@ 6.3.1 Linear Policies
A linear controller $K\in{\mathbb{R}}^{d_{u}\times d_{x}}$ is a linear operator that maps the state to the control. We denote by $\Pi_{\gamma}^{L}$ the set of all linear controllers that are $\gamma$-stabilizing according to definition 2.4 and are bounded in Frobenius norm by $\|K\|\leq\gamma$.
@@ 6.3.2 Disturbance Action Controllers
Any DAC policy in $\Pi_{h,\gamma}^{D}$ is $2\gamma^{2}$-stabilizing.
@@
For convenience, we define a fictitious noise term $\tilde{\mathbf{w}}_{t}=B\sum_{i=1}^{h}M_{i}\mathbf{w}_{t-i}$. First, due to the bound on $M_{1:h}$ and $\mathbf{w}_{t}$, we have that
$\|\tilde{\mathbf{w}}_{t}\|=\left\|\sum_{i=1}^{h}M_{i}\mathbf{w}_{t-i}\right\|\leq\sum_{i=1}^{h}\|M_{i}\|\|\mathbf{w}_{t-i}\|\leq\gamma,$
@@
$=A\mathbf{x}_{t}+B\left(K\mathbf{x}_{t}+\sum_{i=1}^{h}M_{i}\mathbf{w}_{t-i}\right)+\mathbf{w}_{t}$
$=(A+BK)\mathbf{x}_{t}+B\sum_{i=1}^{h}M_{i}\mathbf{w}_{t-i}+\mathbf{w}_{t}$
$=\sum_{j=0}^{t}\left((A+BK)^{j}\left(B\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i-j}+\mathbf{w}_{t-j}\right)\right)$
$=\sum_{j=0}^{t-1}\left((A+BK)^{j}(B\tilde{\mathbf{w}}_{t-j}+\mathbf{w}_{t-j})\right).$
We can thus view the states $\mathbf{x}_{t}$ as generated by a $\gamma$-stabilizing policy over noises that are of magnitude $2\gamma$, and hence bounded by $2\gamma^{2}$. ∎
@@
$\leq\left\|\sum_{i=0}^{t}A^{i}B\right\|\max_{i\leq t}\|\mathbf{u}_{i}-\tilde{\mathbf{u}}_{i}\|\leq\frac{\kappa}{\delta}\varepsilon,$
@@ 6.4 Bibliographic Remarks
Regret minimization, online learning, and online convex optimization have a rich history and have been studied for decades in the machine learning and game theory literature. The reader is referred to the comprehensive texts of [CL06, HAZ16] for a detailed treatment.
Linear state-feedback policies are a mainstay of control theory in both applications, due to their simplicity, and in theory, due to their optimality in deterministic and stochastic settings, as demonstrated in [KAL60]. The methods for computing such policies are often based on dynamic programming. Often motivated by the need to impose constraints beyond optimality, such as robustness, various other policy parameterizations were introduced. An early example is the Youla parameterization, sometimes called Q-parameterization [YJB76]. Later, the introduction of system-level synthesis (SLS) [DMW+17, ADL+19], in a similar vein, proved useful for distributed control, especially given its emphasis on state-space representations. For a comparison of these and additional ones such as input-output parameterizations [FZP+19], see [ZFP+20, ZZL+23].
The Disturbance Action Control parameterization, introduced in [ABH+19b], is similar to system-level synthesis. Like SLS, it is a convex parameterization that is amenable to the tools of convex optimization. However, unlike SLS, it is a more compact representation that has logarithmically many parameters in its natural representation in terms of the desired approximation guarantee against linear policies. This makes it better suited for machine learning.
The study of regret in the context of control has roots in the work of [AS11, ABK14]. These works, as well as many that followed [DMM+18, MTR19, CHK+18, AHS19c, SF20], study regret in a stochastic setting, where the perturbations are stochastic. In such settings, optimal control is well defined. The full power of regret analysis for control, involving both adversarial perturbations, adversarial loss functions, and computational considerations, was initiated in [ABH+19b]. The notion of regret as applied to policy classes was implicit in the work of [EKM04]. The regret metric when applied to policies is different from the counterfactual notion of policy regret defined by [ADT12].
@@ 7.4 Bibliographic Remarks
The bulk of the material in this chapter is from [ABH+19b]. The work of [ABH+19b] defined the general online nonstochastic setting in which the disturbances and the cost functions are adversarially chosen, and where the cost functions are arbitrary convex functions.
The follow-up work by [AHS19c] achieves a logarithmic pseudo-regret for strongly convex, adversarially selected losses, and well-conditioned stochastic noise. Logarithmic regret algorithms for the full setting of online nonstochastic control with strongly convex costs were obtained in [SIM20, FS20, CCK20].
Online nonstochastic control with known dynamics was further extended in several directions. The task of control with bandit feedback for the costs was considered in [GHH20a]. More refined regret metrics applicable to changing dynamics were studied in [GHM20b, MGS+21, ZWZ22, SIN22]. More extensions suitable for changing environments were studied in [BW22]. Online nonstochastic control of population dynamics was studied in [GHL+24].
[LDL21] studies algorithms that balance regret minimization while adhering to constraints. An application of nonstochastic control in iterative planning in the presence of perturbations is described in [AHM+21]. Various open-source software packages, such as [GHM20b], provide benchmarked implementations of many such algorithms.
An alternate performance metric for online control, which is also nonstochastic, is the competitive ratio. This objective tracks the cost of the online controller and compares it to the best policy in hindsight. Control methods that minimize the competitive ratio were initiated in [GW19]. This approach was extended in a series of papers [GLS+19, SLC+20, GH21]. Thus far, the success of this approach was limited to fully observed linear dynamical systems and further to quadratic costs when achieving the optimal competitive ratio. [GAS+22] show that regret minimizers in the non-stochastic setting automatically achieve a near-optimal competitive ratio.
A precursor to online nonstochastic control is a line of work, beginning with [ABK14], focused on the control of stochastic systems under sequentially revealed adversarial costs functions. See [CK20, CCK22] for further developments.
@@ 8.6 Bibliographic Remarks
In contrast, much of the work at the intersection of machine learning and control is focussed on precise non-asymptotic bounds on sample complexity or regret. [FIE97] which studies the discounted LQR problem under stochastic noise is the earliest example that provides a precise sample complexity bound. Similarly, the early work of [AS11] uses an optimism-based approach to achieve optimal regret in terms of dependence on horizon. This regret bound was later improved for sparse systems in [IJR12]. [DMM+18, DMM+20] established regret and sample complexity results for the general stochastic LQR formulation with a polynomial dependence on natural system parameters. [FGK+18] studies the efficacy of first-order optimization-based approaches in synthesizing controllers for the LQR problem. The work of [MTR19] establishes that simple explore-then-commit strategies achieve the optimal dependence of regret on horizon, with [SF20] proving that such approaches are optimal in terms of their dependence on dimension of the system, as well. [CKM19] also gave a first optimal-in-horizon polynomial regret result simultaneously with the former work, albeit using an optimism-based approach. The result of [PH20] gives a rate-optimal regret bound for unknown systems with convex costs.
@@ 9.1 Learning in Dynamical Systems
For the remainder of this chapter, we consider a partially observable linear time-invariant dynamical system described by the following equations.
$\mathbf{x}_{t+1}$ $=A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t},$
@@ 9.2 Online Learning of Dynamical Systems
In the problem of online learning in a linear dynamical system, sequentially for $t=1,\ldots,T$, the online learning algorithm:
  • observes input $\mathbf{u}_{t-1}\in{\mathbb{R}}^{d_{u}}$,
  • predicts output $\hat{\mathbf{y}}_{t}\in{\mathbb{R}}^{d_{y}}$,
  • observes true system output $\mathbf{y}_{t}$,
  • suffers loss according to a pre-specified loss function, $\ell_{t}(\hat{\mathbf{y}}_{t},\mathbf{y}_{t})$, e.g.
    $\ell_{t}(\hat{\mathbf{y}}_{t},\mathbf{y}_{t})=\|\hat{\mathbf{y}}_{t}-\mathbf{y}_{t}\|^{2}.$
The goal of the online learner is to minimize total loss over all game iterations. If the observations ${\mathbf{y}}_{t}$ are arbitrary, this is an ill-defined goal, and rather than minimizing loss, we consider a comparative analogue.
More precisely, we consider a family of prediction rules from available information to prediction, and consider the metric of regret with respect to the best of these rules in hindsight. Formally, let $\Pi$ be a family of mappings from $\{\mathbf{u}_{\tau},\mathbf{y}_{\tau}\}_{\tau=1}^{t-1}$ to a prediction $\hat{\mathbf{y}}_{t}$, i.e.
$\pi\in\Pi\ ,\ \pi(\mathbf{u}_{1:t-1},\mathbf{y}_{1:t-1})=\hat{\mathbf{y}}_{t}^{\pi}\in{\mathbb{R}}^{d_{y}}.$
Let $\mathcal{A}$} be a prediction algorithm that iteratively produces predictions $\hat{\mathbf{y}}_{t}$. Then its regret with respect to $\Pi$ is defined as
$\mathrm{{regret}}_{T}({\mathcal{A}},\Pi)=\sum_{t=1}^{T}\ell_{t}(\hat{\mathbf{y}}_{t},\mathbf{y}_{t})-\min_{\pi\in\Pi}\sum_{t=1}^{T}\ell_{t}(\hat{\mathbf{y}}^{\pi}_{t},\mathbf{y}_{t}).$
We proceed to define common prediction classes in the context of linear dynamical systems.
@@ 9.3 Classes of Prediction Rules
We say that a class of predictors $\Pi_{1}$ $\varepsilon$-approximates class $\Pi_{2}$ over horizon $T$ if the following holds. For every sequence of inputs and outputs $\{\mathbf{u}_{1:T},\mathbf{y}_{1:T}\}$ and every $\pi_{2}\in\Pi_{2}$, there exists $\pi_{1}\in\Pi_{1}$ such that
$\sum_{t=1}^{T}\|\hat{\mathbf{y}}_{t}^{\pi_{1}}-\hat{\mathbf{y}}_{t}^{\pi_{2}}\|\leq T\varepsilon,$
where
$\hat{\mathbf{y}}_{t}^{\pi_{1}}=\pi_{1}(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1})\ ,\ \hat{\mathbf{y}}_{t}^{\pi_{2}}=\pi_{2}(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}).$
@@
Assume that each loss function $\ell_{t}(\hat{\mathbf{y}},\mathbf{y}_{t})$ is $G$-Lipschitz in its first argument on the relevant prediction domain. If predictor class $\Pi_{1}$ $\varepsilon$-approximates class $\Pi_{2}$, then for any prediction algorithm $\mathcal{A}$} we have
$\mathrm{{regret}}_{T}({\mathcal{A}},\Pi_{2})\leq\mathrm{{regret}}_{T}({\mathcal{A}},\Pi_{1})+G\varepsilon T.$
@@ 9.3.1 Linear Predictors
The simplest prediction rule is a linear function of past observations and past inputs.
Definition 9.3 (Linear Predictors).
A linear predictor parameterized by matrices
$M_{1,i}\in{\mathbb{R}}^{d_{y}\times d_{y}}\ \text{ for }\ i=1,\ldots,h,\qquad M_{2,j}\in{\mathbb{R}}^{d_{y}\times d_{u}}\ \text{ for }\ j=1,\ldots,k,$
maps past outputs and inputs to a prediction according to
@@
We denote by $\Pi_{h,k,\kappa}^{L}$ the set of all such predictors satisfying
$\sum_{i=1}^{h}\|M_{1,i}\|+\sum_{j=1}^{k}\|M_{2,j}\|\leq\kappa.$
This class is computationally attractive because it is parameterized by a convex set of matrices. At first sight, however, it may appear too restrictive to model the predictions generated by a linear dynamical system with hidden state. The next subsection explains why this concern is misleading.
@@ 9.3.2 Linear Dynamical Predictors
An intuitive class of prediction rules that makes use of the special structure of linear dynamical systems is that of open-loop linear dynamical predictors.
The motivation for this class is the case that the system transformations are known and the system is fully observed, namely $C=I$, the identity matrix. In this situation, the optimal prediction can be calculated by unrolling the system evolution equation $\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t}$. It can be seen that this prediction rule is optimal for any zero-mean noise or perturbation model. However, this optimality is only true if the state is fully observable. In the next chapter, we consider optimal predictors for zero-mean i.i.d. noise in the more general partially observed case.
@@
$=CA^{t}\mathbf{x}_{0}+\sum_{i=1}^{t}CA^{i-1}B\mathbf{u}_{t-i}.$
Thus, the operator $\pi_{\{A,B,C,\mathbf{x}_{0}\}}$ takes a history of inputs and gives the next output of the system accurately. We denote by $\Pi^{\star}$ the set of all such predictors that in the case of zero noise give a perfectly accurate prediction.
The class $\Pi^{\star}$ is a very natural comparator class for the noiseless open-loop prediction problem. It is important to note, however, that in the general partially observed noisy setting the optimal predictor need not belong to this class, since it may depend on past outputs in a nontrivial way. We return to this more general question in the next chapter. For the present chapter, $\Pi^{\star}$ serves as a clean structural benchmark whose induced predictors we can approximate or represent by convex classes.
9.4 Convex Relaxations of Linear Dynamical Predictors
In the previous sections we have defined what it means to learn in a dynamical system, and described two natural classes of predictors. The class of linear dynamical predictors is most natural, as it exactly captures the noiseless evolution of a linear dynamical system. However, it is nonconvex in its parametrization.
In this section we describe convex relaxation of this class that approximate it extremely well and allow for efficient learning.
9.4.1 Convex Relaxation of Stable Systems
The first relaxation we discuss is that which assumes additional structure from the dynamics, namely stability. For a noiseless system, define the truncated open-loop predictor
$\hat{\mathbf{y}}_{t}=\pi_{\{A,B,C,h\}}(t)=\sum_{i=1}^{h}CA^{i-1}B\mathbf{u}_{t-i},$
where by convention $\mathbf{u}_{s}=0$ for all $s\leq 0$. Denote by $\Pi_{h}^{\star}$ the class of all such truncated predictors.
When the system is stable, this truncation yields a simple approximation to the class of open-loop linear dynamical predictors. To state this in the language of Definition 9.1, we restrict attention to a stable subclass with uniformly bounded parameters and zero initial state.